# The Role of Inclusion, Control, and Ownership in Workplace AI-Mediated Communication

### v1

@@ 4.2. Control
Figure 3 illustrates, for each gender condition (different rows) the frequency of different survey responses for the following questions: how much control did you feel over the process of writing the message (“Control During Process” on the left), and how much control did you feel over the final version of the message (“Control Over Message” on the right). Figure 3a (top-left) shows in the condition with the self-assured AI writing style suggestions, during the writing process, 48% of men and 49% of WNBSD participants reported feeling a lot or great deal of control (indicated by the darker blue and darkest blue bars). In the condition with the hesitant AI writing style, seen in Figure 3c (bottom-left), 61.1% of the men and 61.2% of the WNBSD participants expressed a lot or a great deal of control during the writing process. A two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05), meaning that participants in the hesitant condition reported more control over the process. There were no significant differences for gender, or interaction between gender and AI treatment group for this question.
In both treatment conditions, participants generally reported higher levels of control over the final version of the message than the writing process. In the condition with the self-assured AI writing style, (Figure 3b, top-right), 59.6% of men and 70.6% of WNBSD participants reported a lot or a great deal of control during the writing process. Figure 3d (bottom-right) shows 73% of men and 80.2% of WNBSD participants who wrote with a hesitant writing style model reported a lot or a great deal of control over the final version of the message. For this question as well, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05), with participants in the hesitant AI writing style condition reported greater control over the final version of the message. At the same time, there were no significant differences in relation to gender or interactions between gender and AI treatment group for this question.
@@ 4.3. Ownership
Figure 4 illustrates the frequency of different survey responses for each gender condition (rows) for the following questions: to what extent do you feel like the message you wrote is yours (“Message Is Yours”), and thinking back on the message writing activity, how much did the message sound like you (“Message Sounds Like You”). In the self-assured AI writing style condition, (Figure 4a, top-left), 41.4% of men and 44.1% of WNBSD participants expressed a lot or a great deal of ownership over the message (as seen by the dark blue and darkest blue bars). In the hesitant AI writing style condition (Figure 4c, bottom-left), participants experienced a higher level of ownership with 55.6% of the men and 53.5% of the WNBSD participants expressing a similar degree of ownership over the message. A two-way ANOVA revealed that this difference between the AI treatment groups was significant (p¡0.05). There was no significant difference for gender, or interaction between gender and AI treatment group for this question.
Reflecting on the style of the message, in the self-assured AI writing style condition figures 4b and d on the right show similar trends: participants in the hesitant AI writing style condition (bottom) were more likely to say the final message reflected their style. Once again, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p¡0.05). There was no significant difference for gender, or interaction between gender and AI treatment group and gender for this question.
Overall, our results show AI writing style impacts control and ownership. Participants who wrote with the hesitant model, expressed greater control over the writing process and the final message and felt the message sounded more like them than participants who wrote with the self-assured model. However, our analysis did not reveal an impact of the AI writing style on perceived inclusion. Additionally, we did not see any effects of gender or interaction between gender and treatment group across inclusion, control, and ownership questions. We further explore these factors in the analysis below.
@@ 4.4. Exploratory Analysis: Inclusion and Agency
To better understand the relationship between the different measures and the participants’ interaction with the AI model, we perform an exploratory analysis that brings together multiple factors using a linear regression model. To perform this analysis, we combined various measures into simplified high-level constructs. First, we use as our dependent variable a compound measure of agency. The concept of agency has been defined in several different ways, but broadly captures the idea of “the capacity to alter a situation” (Neff and Nagy, 2018; Cesafsky et al., 2019). The variables that reflect that in our study were control and ownership. While distinct, the literature shows these concepts are closely related: whether a person has control greatly influences their perceived ownership (Mieczkowski, 2022; Braun et al., 2018). An analysis of our measures confirmed the control and ownership questions were indeed highly correlated in our data. We thus created a score for agency by performing a row-wise average across the control and ownership questions.
The independent variables in the model consisted of a simplified inclusion variable, the AI reliance score, the AI writing style, and demographic variables. To simplify the model and analysis, we combined the two highly correlated inclusion measures into one variable by performing a row-wise average across the two inclusion questions. The model also included the measure of AI Reliance as defined above, namely how much the participants accepted the AI suggestions in the writing process. Finally, we include in our model the AI writing style of the model that participants wrote with (the AI treatment), as well as more detailed demographic variables: age, gender, race, and income. The race and gender features were operationalized as binary variables where 1 reflected the privileged group and 0 reflected the minoritized group (e.g., for gender, 1=men, 0=WNBSD). We also mapped the AI writing style as a binary variable (0=hesitant model, 1=self-assured model). We operationalized the age and income demographic features as ordinal variables. Note again that we consider this analysis exploratory, as it was not part of the preregistered analysis presented above that focused on the main effects.
@@
We explore the interaction effects to better understand their impact on perceived agency. Figure 4(a) depicts the interaction between AI reliance and perceived inclusion. AI reliance is the x-axis and perceived agency is the y-axis. The graph compares the groups that showed high (dark blue), medium (blue), and low (light blue) levels of inclusion. Participants with a high level of inclusion, as indicated by the dark blue line in the figure, maintain their sense of agency even with increasing reliance on the writing assistant. Participants with a low (and to a lesser degree, medium) sense of inclusion lose their sense of agency rapidly when they rely more on AI.
@@ 5. Discussion
Our findings enhance the current models of “writing with AI” by introducing new factors and furthering our understanding of their relationships. The findings also hint at the role that gender may play in such context, though this line of investigation requires additional substantiation. Overall, the findings offer implications for research and design of co-writing interactions with AI.
@@
By introducing the construct of inclusion, our work extends the previous research on AI-MC which has examined perceptions of agency while interacting with an AI writing assistant (Lee et al., 2022; Mieczkowski, 2022; Mieczkowski et al., 2021) We measure the concept of inclusion after the co-writing process, by asking participants if they could envision themselves in the model and whether the model’s writing style matched their own. We show that this concept contributes to the perception of agency, a combined metric reflecting ownership and control over the writing process and the final product.
@@
Informed by these findings, Figure 6 presents a new conceptual model of the factors contributing to agency in AI co-writing environments. This model is tentative: not all of these concepts were measured, let alone tested, in our work here. The model suggests that there are two paths for feeling agency in the co-writing process and its outcome. On the top, AI writing style affects AI-user alignment—which we define as the similarity between the AI writing style and the user’s own preferences, that may be derived from their identity. High AI-user alignment leads to increased perceived inclusion that will positively contribute to perceived agency. In our research, we did not measure AI-user alignment. Instead, we built on the literature suggesting that minoritized genders are less likely to be aligned with assertive AI suggestions – which was not the case in this study. At the bottom of the model, AI writing style affects AI-task alignment, which we define as the perceived usefulness of the AI assistant for the task (e.g. writing to ask for a promotion with an assertive AI). If there is high AI-task alignment, the user will rely on the model to complete the task. However, such reliance, as we have shown, can negatively contribute to the user’s perceived agency. The conceptual model includes an interaction between AI reliance and inclusion, as suggested by our exploratory results.
@@
Our model raises several open questions that have implications for designers. The most salient question is the potential competition between AI-User alignment and AI-Task alignment. Or, how can developers strike a balance between models that are inclusive of various users but are also effective in completing the task? We note that even if the models align with both user and task, there is some conflict: the AI-User alignment will ultimately contribute to agency, but the AI alignment with the task, which will increase reliance, will reduce agency. This finding complicates the question of how developers can build models that help the user accomplish the task while also supporting the user’s agency, when there is tension between those.
@@
While gender differences did not account for our expected main effect, our exploratory analysis also suggests there are still, potentially, some differences between genders. A noteworthy finding in this analysis was the significant interaction between gender and inclusion in relation to perceived agency. The interaction suggests that for minoritized genders, perceptions of inclusion have a greater effect on the perception of agency compared to men. If they hold, these findings lend support to our initial hypothesis: that people of different genders may react differently to writing assistants of different styles. This finding suggests that it may be useful to devise, develop, and test, language models that also optimize on generating perceptions of inclusion by people of different demographic backgrounds.

### v2

@@ 4.2. Control
Figure 3 illustrates the frequency of different survey responses for each gender (different columns) for the following questions: how much control did you feel over the process of writing the message (“Control During Process” on the left), and how much control did you feel over the final version of the message (“Control Over Message” on the right). Figure 3a shows that in the condition with the hesitant AI writing style suggestions (left half), during the writing process, 61.1% of the men (first column) and 61.2% of WNBSD participants (second column) reported feeling a lot or great deal of control (the darker blue and darkest blue bars in the columns). In the condition with the self-assured AI writing style (right half), 48% of men and 49% of the WNBSD participants expressed a lot or a great deal of control during the writing process. A two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.0044), meaning that participants in the hesitant condition reported more control over the process. There were no significant differences for gender or the interaction between gender and the AI treatment groups for this question. In both treatment conditions, participants generally reported higher levels of control over the final version of the message than the writing process.
In the condition with the hesitant AI writing style (Figure 3b, left), 73% of men and 80.2% of WNBSD participants reported a lot or a great deal of control over the final version of the message. Figure 3b (right) that shows 59.6% of men and 70.7% of WNBSD participants who wrote with a self-assured writing style model reported a lot or a great deal of control over the final version of the message. For this question as well, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.046), with participants in the hesitant AI writing style condition reporting greater control over the final version of the message. At the same time, there were no significant differences in relation to gender or interactions between gender and the AI treatment group for this question.
@@ 4.3. Ownership
Figure 4 illustrates the frequency of different survey responses for each gender condition (columns) for the following questions: to what extent do you feel like the message you wrote is yours (“Message Is Yours”), and thinking back on the message writing activity, how much did the message sound like you (“Message Sounds Like You”). In the hesitant AI writing style condition (Figure 4a, left), 55.6% of the men and 53.5% of WNBSD participants expressed a lot or a great deal of ownership over the message (the dark blue and darkest blue bars in the column). In the self-assured AI writing style condition (Figure 4a, right), participants experienced a higher level of ownership, with 41.4% of men and 44.1% of the WNBSD participants expressing a similar degree of ownership over the message. A two-way ANOVA revealed that this difference between the AI treatment groups was significant (p=). There was no significant difference for gender or the interaction between gender and the AI treatment groups for this question.
Reflecting on the style of the message, Figure 4b shows a similar trend: participants in the hesitant AI writing style condition (left) were more likely to say the final message reflected their style. Once again, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.032). There was no significant difference for gender or the interaction between gender and the AI treatment groups and gender for this question.
Overall, our results show AI writing style impacts control and ownership. Participants who wrote with the hesitant model expressed greater control over the writing process and the final message and felt the message sounded more like them than participants who wrote with the self-assured model. However, our analysis did not reveal an impact of the AI writing style on perceived inclusion. Additionally, we did not see any effects of gender or interaction between gender and treatment group across inclusion, control, and ownership questions. We further explore these factors in the analysis below.
@@ 4.4. Exploratory Analysis: Inclusion and Agency
To better understand the relationship between the different measures and the participants’ interaction with the AI model, we perform an exploratory analysis that brings together multiple factors using a linear regression model. To perform this analysis, we combined various measures into simplified high-level constructs. First, we use a compound measure of agency as our dependent variable. The concept of agency has been defined in several different ways but broadly captures the idea of “the capacity to alter a situation” (Neff and Nagy, 2018; Cesafsky et al., 2019). The variables that reflect that in our study were control and ownership. While distinct, the literature shows these concepts are closely related: whether a person has control greatly influences their perceived ownership (Mieczkowski, 2022; Braun et al., 2018). An analysis of our measures confirmed the control and ownership questions were indeed highly correlated in our data. We thus created a score for agency by performing a row-wise average across the control and ownership questions.
The independent variables in the model consisted of a simplified inclusion variable, the AI reliance score, the AI writing style, and demographic variables. To simplify the model and analysis, we combined the two highly correlated inclusion measures into one variable by performing a row-wise average across the two inclusion questions (Table 3 in the appendix shows the correlation of the inclusion, control, and ownership measures). The model also included the measure of AI reliance as defined above, namely how much the participants accepted the AI suggestions in the writing process. Finally, we include in our model the AI writing style of the model that participants wrote with (the AI treatment), as well as more detailed demographic variables: age, gender, race, and income. The race and gender features were operationalized as binary variables where 1 reflected the privileged group and 0 reflected the minoritized group (e.g., for gender, 1=men, 0=WNBSD). We also mapped the AI writing style as a binary variable (0=hesitant model, 1=self-assured model). We operationalized the age and income demographic features as ordinal variables. Note again that we consider this analysis exploratory, as it was not part of the preregistered analysis presented above that focused on the main effects.
@@
We explore the interaction effects to better understand their impact on perceived agency. Figure 4(a) depicts the interaction between AI reliance and perceived inclusion. AI reliance is the x-axis, and perceived agency is the y-axis. The graph compares the groups that showed high (dark blue), medium (blue), and low (light blue) levels of inclusion. Participants with a high level of inclusion, as indicated by the dark blue line in the figure, maintain their sense of agency even with increasing reliance on the writing assistant. Participants with a low (and to a lesser degree, medium) sense of inclusion lose their sense of agency rapidly when they rely more on AI.
@@ 5. Discussion
Our findings enhance the current models of “writing with AI” by introducing new factors and furthering our understanding of their relationships. The findings also hint at the role that gender may play in such a context, though this line of investigation requires additional substantiation. Overall, the findings offer implications for research and design of co-writing interactions with AI.
@@
By introducing the construct of inclusion, our work extends the previous research on AI-MC, which has examined perceptions of agency while interacting with an AI writing assistant (Lee et al., 2022; Mieczkowski, 2022; Mieczkowski et al., 2021). We measure the concept of inclusion after the co-writing process by asking participants if they could envision themselves in the model and whether the model’s writing style matched their own. We show that this concept contributes to the perception of agency, a combined metric reflecting ownership and control over the writing process and the final product.
@@
Informed by these findings, Figure 6 presents a new conceptual model of the factors contributing to agency in AI co-writing environments. This model is tentative: not all of these concepts were measured, let alone tested, in our work here. The model suggests that there are two paths for feeling agency in the co-writing process and its outcome. On the top, AI writing style affects AI-user alignment—which we define as the similarity between the AI writing style and the user’s own preferences that may be derived from their identity. High AI-user alignment leads to increased perceived inclusion, positively contributing to perceived agency. In our research, we did not measure AI-user alignment. Instead, we built on the literature suggesting that minoritized genders are less likely to be aligned with assertive AI suggestions – which was not the case in this study. At the bottom of the model, AI writing style affects AI-task alignment, which we define as the perceived usefulness of the AI assistant for the task (e.g., writing to ask for a promotion with an assertive AI). If there is high AI-task alignment, the user will rely on the model to complete the task. However, such reliance, as we have shown, can negatively contribute to the user’s perceived agency. The conceptual model includes an interaction between AI reliance and inclusion, as suggested by our exploratory results.
@@
Our model raises several open questions that have implications for designers. The most salient question is the potential competition between AI-User alignment and AI-Task alignment. Or, how can developers strike a balance between models that are inclusive of various users but are also effective in completing the task? We note that even if the models align with both the user and task, there is some conflict: the AI-User alignment will ultimately contribute to agency, but the AI alignment with the task, which will increase reliance, will reduce agency. This finding complicates the question of how developers can build models that help the user accomplish the task while also supporting the user’s agency when there is tension between those. Further work is needed to understand the trade-off between agency and task performance in real-world applications.
@@
While gender differences did not account for our expected main effect, our exploratory analysis also suggests there are still, potentially, some differences between genders. A noteworthy finding in this analysis was the significant interaction between gender and inclusion in relation to perceived agency. The interaction suggests that for minoritized genders, perceptions of inclusion have a greater effect on the perception of agency compared to men. If they hold, these findings lend support to our initial hypothesis: that people of different genders may react differently to writing assistants of different styles. This finding suggests that it may be useful to devise, develop, and test language models that also optimize on generating perceptions of inclusion by people of different demographic backgrounds.
Our work explored the impact of AI auto-complete suggestions on feelings of inclusion, control, and ownership over the written text and the implications for users of various backgrounds. The work has a number of limitations. First, in asking people to write as if asking for a promotion, our study focused on an important and realistic workplace task. However, other workplace writing tasks should also be studied to validate the results. Furthermore, we did not measure individual differences in assertiveness, which could have provided additional insight on the effect of the treatment. Pre-task measurement of such attitudes, though, could have introduced bias to the experimental task. We focus on English-speaking US residents. Additional work is needed to understand how users from various linguistic backgrounds and cultures interact with such AI agents. Despite these limitations, this short paper showed some of the first evidence that minoritized genders’ perceptions of control and ownership could vary in some ways from men’s as a response to writer-style biases. We offer a new conceptual model for understanding the impact of these factors on people’s feelings of overall agency in AI-mediated environments.
Appendix A Appendix
A.1. Validating the Treatment
We used human annotators and computational methods to validate the effect of the auto-complete suggestions. We are able to analyze the auto-complete suggestions as well as the final message, as our writing app logs all auto-complete suggestions seen by users. For our analysis, we randomly sampled 100 participants from each AI treatment condition. We selected suggestions that appeared at the beginning, middle, and end of the writing task for a total of five suggestions per participant. We used LIWC-22 on the auto-complete suggestions to analyze the following features: positive tone, negative tone, pro-social behavior, politeness, conflict, clout, authenticity, and social behavior as a whole. Clout refers to the social status or confidence people display in writing. Authenticity refers to how much a person self-regulates, with low authenticity indicating the person is socially cautious and high authenticity indicating a person is spontaneous. Based on prior literature (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008), self-assured language and hesitant language should differ along these metrics. With the exception of the positive tone metric, t-tests revealed significant differences between the means across all other the metrics (negative tone: t=-6.73, p=; pro-social behavior: t=3.29, p=0.0012; politeness: t= 3.77, p=0.0024, conflict: t=-2.27, p=0.025; clout: t=3.15, p=0.002; authenticity: t=-4.00, p=; social behavior: t=3.24, p=0.001). Table 2 shows several of the suggestions with some selected self-assured suggestions highlighted in dark blue and selected hesitant suggestions highlighted in light blue.
| Prompt | Sampled Suggestions |
Come up with a continuation to the input text in a self-assured manner asking for a raise |
“Be justifiable to review my current compensation package. I believe an increase commensurate with my increased…” “Over the past year, I have consistently met and exceeded my performance objectives, taking on additional responsibilities without…” “My compensation should reflect this. I would appreciate the opportunity to discuss this with you.” “It is fair to align my compensation with these mounting responsibilities. I have done extensive research on industry.” |
Come up with a continuation to the input text in a hesitant manner asking for a raise |
“I certainly don’t want to come across as ungrateful or discontented.” “…my performance and contributions, if needed. I am not asking for an immediate response, but rather a thoughtful…” “In comparison to the industry standard, my current compensation seems a bit… well, it’s not entirely…” “I thought it might be time to discuss the possibility of… well, perhaps a slight adjustment in my…” |
Additional analysis showed the suggestions had an effect on the final message. We had three human annotators evaluate all of the final messages on a scale from 1-10, where 1 indicated no assertiveness and 10 indicated high assertiveness. T-tests revealed the means between conditions to be significant (self-assured/control: t=14.56, p=; self-assured/hesitant: t= 14.56, p=; control/hesitant: t= 10.09, p=). For the LIWC-22 analysis, t-tests showed significant differences between groups, with the exception of the positive tone metric and politeness (negative tone: t=-7.09, p=; pro-social behavior: t=4.83, p=; conflict: t=-2.24, p=0.026; clout: t= 3.00, p=0.003; authenticity: t= -2.24, p=0.03; social behavior: t= 5.36, p=).
A.2. Exploratory Analysis Details
In the exploratory analysis (section 4.4), we performed a row-wise average across the inclusion questions and the control and ownership questions to create our agency measure. Table 3 shows the correlation between the post-task questions. The two ownership questions (message sounds like me, message is mine), two control questions (control during process, control over message), and two inclusion questions (assistant made for me, assistant sounds like me) are highly correlated, as indicated by the bold text.
| message is mine | control over message | control during process | message sounds like me | assistant made for me | (assistant sounds like me) | ai_reliance | |
| message is mine | 1.00 | ||||||
| control over message | 0.53 | 1.00 | |||||
| control during process | 0.45 | 0.65 | 1.00 | ||||
| message sounds like me | 0.67 | 0.51 | 0.45 | 1.00 | |||
| assistant made for me | 0.23 | 0.21 | 0.21 | 0.37 | 1.00 | ||
| assistant sounds like me | 0.13 | 0.17 | 0.22 | 0.21 | 0.56 | 1.00 | |
| ai_reliance | -0.37 | -0.27 | -0.22 | -0.28 | 0.13 | 0.16 | 1.00 |
Table 4 shows the correlation analysis for the predictors in our model (Table 1 in section 4.4). The demographic features race and gender in the model are binary with 1 corresponding to the privileged group (white, male) and 0 corresponds to minoritized groups (non-white, WNBSD).
| Age | Race | Income | Gender | AI Reliance | Inclusion | AI Style | |
| Age | 1.00 | ||||||
| Race | 0.16 | 1.00 | |||||
| Income | 0.22 | 0.06 | 1.00 | ||||
| Gender | -0.02 | -0.05 | 0.08 | 1.00 | |||
| AI Reliance | 0.04 | -0.01 | 0.09 | 0.07 | 1.00 | ||
| Inclusion | -0.08 | -0.12 | -0.08 | 0.02 | 0.17 | 1.00 | |
| AI Style | -0.00 | 0.03 | 0.04 | -0.10 | 0.14 | 0.10 | 1.00 |